Main INDEX
Monthly INDEX
PREV
NEXT
User name gilman
Log entry time 11:09:15 on September21,2006
Entry number 184338
keyword=target problem
The target was nice and stable, until for reasons I do not understand - presumably the 15 k cryoflows decreased with the Hall C cooldown? - the targets warmed up to 21.1 K for LH2 and 23.4 K for LD2. m Seely came by and suggested our flow was low so we started opening the JT valves. It appears that the deuterium JT valve closed slowly when stepped open, as the deuterium slowly got warmer and the valve position readback appeared to decrease. M Seely suggested a reboot of the ioc, telling us that they had a problem before in G0 where the direction bit in the controller appeared to latch in the close direction, so steps in either direction closed the JT valve.
The staus right now is, after rebooting the ioc, the JT valves are at about 34.7 (LH2) and 15.5 (LD2), and the target temperatures are about 19.5 and 27.0 K and cooling slowly back down. The total 15 K flow is about 10 g/s; M Seely indicated we should have no problems running at 10 g/s flow - in terms of exceeding our allocation of the available coolant.